Chris bought a $59,500 piece of equipment for his private company. The payment plan called for payments of $7,000 every half-year for 7 years. The effective annual interest rate charged to Chris is closest to … Note: The first payment occurred 6 months after the purchase.

a) 5.51%

b) 7.49%

c) 8.50%

d) 11.37%

e) 14.98%

f) 15.54%

g) 17.30%
Answer #1

Time is 14 half years and periodic payments are 7000. This gives a cash flow equation

-59500 + 7000(P/A, i%, 14) = 0

For i = 7% NPV = 1718.28

For i = 8% NPV = - 1790.34

This gives i = 7% + (8% - 7%)*(1718.28/(1718.28+1790.34))

= 7.49%

Effective annual rate = (1 + 7.49%)^2 = 15.54%

Select option F.